### v1

# Label-Looping: Highly Efficient Decoding for Transducers

Abstract
This paper introduces a highly efficient greedy decoding algorithm for Transducer inference. We propose a novel data structure using CUDA tensors to represent partial hypotheses in a batch that supports parallelized hypothesis manipulations. During decoding, our algorithm maximizes GPU parallelism by adopting a nested-loop design, where the inner loop consumes all blank predictions, while non-blank predictions are handled in the outer loop. Our algorithm is general-purpose and can work with both conventional Transducers and Token-and-Duration Transducers. Experiments show that the label-looping algorithm can bring a speedup up to 2.0X compared to conventional batched decoding algorithms when using batch size 32, and can be combined with other compiler or GPU call-related techniques to bring more speedup. We will open-source our implementation to benefit the research community. Equal contribution.
keywords:
speech recognition, Transducer, TDT, Token-and-Duration Transducer, parallel computing
1 Introduction
Speech recognition is the task of converting human speech audio into the text sequence that it represents. As speech is one of the most natural ways people communicate, speech recognition technology has been a hot research topic for decades. Nowadays, because of the research efforts and the abundance of computational resources available, we have developed sophisticated speech models that can achieve high accuracy in recognition tasks, and most of those models are freely available with many high-quality open-source speech recognition toolkits, including NeMo [1], ESPNet [2], SpeechBrain [3], Espresso [4], etc.
One popular model for speech processing is Transducer [5]. In terms of model architecture, Transducer and CTC [6] are related in that they both adopt a frame-synchronous paradigm and have a special blank symbol as an output label so that the model output can be of a different length than the audio input. The major difference between those models is that CTC adopts a conditional independence assumption in predicting tokens from each frame. At the same time, Transducers consider textual context during token prediction, making it achieve better accuracy than CTC models at a slightly increased computational cost.
Many research works have attempted to improve the efficiency of Transducer models. More powerful encoder architectures, e.g. Transformers [7] and Conformers [8] are proposed to replace the original LSTM encoders [5], which both improve the model’s accuracy and efficiency. [9] replaced LSTM decoders with a stateless decoder, which shows a speed-up of running Transducer models with a slight degradation in model performance. Multi-blank Transducers [10] introduced big blank tokens that allow skipping of multiple frames during decoding when blank symbols are predicted during inference, bringing significant speed-up and also slightly improved model accuracy; Token-and-Duration Transducer (TDT) [11] models extend the multi-blank model by decoupling the prediction of tokens and durations so that the model can skip multiple frames after each prediction, regardless of whether it’s a blank, bringing even more speed-up to Transducer inference. [12] added additional constraints to Transducers in terms of the maximum labels emitted per frame which shows speed-up for its inference. [13] proposed to use a CTC model to help predict blank symbols in advance, so the overall decoding time of the model can be reduced. [14, 15, 16] all proposed alternative ways to expand search paths to speed up beam search for Transducers.
In this paper, we focus on improving the decoding algorithm of Transducer models by proposing a label-looping algorithm for batched inference, which achieves significant speed-up. The contribution of this paper is as follows,
  1. 1.
    We propose a novel data structure to represent partial hypotheses within each batch, that can be efficiently manipulated using PyTorch [17] CUDA tensor operations.
  2. 2.
    We propose a novel strategy to precompute the encoder and decoder projections before the joiner computations.
  3. 3.
    We propose a novel label-looping algorithm that separates the processing of blank and non-blank emissions of Transducer decoding, maximizing parallelism.
  4. 4.
    Our algorithm brings up to 2.0X speedup compared to existing batched decoding algorithms, and can bring up to 3.2X speedup when combined with technique in compilation and GPU call optimization.
  5. 5.
    Our implementation will be open-sourced through NeMo toolkit [1] so that the results can be easily reproduced.
2 Transducer Backgrounds
The Transducer model is a popular end-to-end speech recognition model architecture, first proposed in [5]. As shown in Figure 1, it consists of an encoder, a decoder, and a joiner. The encoder and decoder extract higher-level information from the audio and the text history context, respectively, and the joiner combines the outputs of the encoder and decoder and computes a probability distribution over the vocabulary. The Transducer model has a special blank symbol in its vocabulary, which serves as outputs for input frames that do not contribute additional labels to the output.
Figure 1: Transducer Architecture
The inference of the Transducer model requires traversing the encoder output frame by frame, and iteratively predicting tokens based on the prediction history. Algorithm 1 shows the inference procedure of the Transducer model on a single audio utterance, where the blank symbol is represented as $\O$. We point out several observations of the algorithm, from lines 11 to 15,
  • $t$ is incremented only after blank predictions and remains unchanged for non-blank predictions.
  • only non-blank predictions are added to the output, and used to update decoder states.
Algorithm 1 Inference of Transducer
1:  input: acoustic input $x$
2:  enc = encoder(x) # output dim is [T, dim]
3:  hyp, state, $t$ = [], decoder.init_state(), 0
4:  while $t<$do
5:     if hyp is [] then
6:        dec, new-state = decoder(state, BOS)
7:     else
8:        dec, new-state = decoder(state, hyp[-1])
9:     token-probs = joiner(enc[$t$], dec)
10:     prediction = argmax(token-probs)
11:     if prediction is not Ø then
12:        hyp.append(prediction)
13:        state = new-state
14:     else
15:        $t$ = $t$ + 1
16:  return hyp
The difference in processing blank and non-blank predictions presents difficulties for efficient batch inference for Transducers, since the time index $t$ for different utterances can increment at different times, and partial hypotheses for different utterances might grow asynchronously as well. This asynchrony makes it hard to implement an efficient batched inference algorithm that can leverage parallel computing fully. A common implementation of batched inference for Transducers is shown in Algorithm 2 (for simplicity, this Algorithm assumes all audio in the batch has the same length $T$; in practice, we need extra checks in the code for individual audio lengths). This algorithm has the following limitations in terms of efficiency:
  1. 1.
    The algorithm increments the time stamp $t$ for different utterances synchronously, and if one utterance predicts a non-blank symbol, all other utterances in the same batch must wait before the whole batch advances to the next time step.
  2. 2.
    Lines 7 and 15 selectively update the decoder states and outputs, depending on whether the last prediction is blank. This means some of the decoder state computation is wasted.
  3. 3.
    At lines 11 to 13, processing of different hypotheses in the batch is done sequentially with a for-loop to handle blank and non-blank predictions differently, which can be time-consuming.
Algorithm 2 Batched Inference of Transducer
1:  input: acoustic input $x_{1},x_{2},...,x_{B}$
2:  encs = encoder($x$) # output dim is [B, T, dim]
3:  hyps, states, t = [[] * B], [decoder.init_state() * B], 0
4:  predictions = [BOS * B]
5:  while $t<$do
6:     d, s = decoder(states, predictions)
7:     decs[i], states[i] = d[i], s[i] where predictions[i] $\neq$ Ø
8:     token-probs = joiner(encs[$t$], decs)
9:     predictions = argmax(token-probs)
10:     while predictions is not all Ø do
11:        for i = 1 to B do
12:           if predictions[i] is not Ø then
13:              hyps[i].append(predictions[i])
14:        d, s = decoder(states, predictions)
15:        decs[i], states[i] = d[i], s[i] where predictions[i] $\neq$ Ø
16:        token-probs = joiner(encs[$t$], decs)
17:        predictions = argmax(token-probs)
18:     $t=t+1$
19:  return hyps
3 Methods
In this Section, we describe our algorithm for batched Transducer inference, which consists of two novel components,
  1. 1.
    a data structure to represent hypotheses in a batch that supports parallelized manipulation of hypotheses and,
  2. 2.
    an algorithm to run inference on Transducers that maximizes the parallelism.
3.1 Representation of batched hypotheses
To represent partial hypotheses during inference, instead of having a single Hypothesis class that stores the tokens, time-stamps, scores, etc, we store information about hypotheses in a batch in one BatchedHyps class, in which all information is stored with CUDA tensors. Operations like adding tokens to partial hypotheses, updating scores and time-stamps etc, are implemented with masked CUDA tensor operations for both correctness and high efficiency. Our design stores the texts of partial hypotheses with a 2D tensor of [B, max-length], where B is the number of utterances in the batch, and the max-length is initialized with a smaller value proportional to the length of the audio and will double the size if this max-length is exceeded. We show a simple pseudo-code for the class below and would advise readers to check our open-sourced implementation for details.
class BatchedHyps:
current_lengths: LongTensor[B]
transcripts: LongTensor[B, max_length]
def add_results(self, add_mask, labels):
self.transcripts[range(B),
self.current_lengths] = labels
self.current_lengths += add_mask
3.2 Label-looping algorithm for conventional Transducers
Our proposed algorithm is shown in Algorithm 3. It is based on the observation that, only non-blank predictions require decoder state updates, so we can maximize the parallelism of computation by having the outer loop (line 6) of the algorithm process only non-blank predictions. The less expensive blank predictions are processed in the inner loop (line 12). With this design, after the decoder is run (line 7), there is no need to selectively keep the updated states, a process that is required in the original Algorithm 2 (lines 7 and 15).
Algorithm 3 Label-looping Algorithm
1:  input: acoustic input $x_{1},x_{2},...,x_{B}$
2:  encs = encoder($x$) # output dim is [B, T, dim]
3:  hyps, state = [[] * B], [decoder.init_state() * B]
4:  predictions = [BOS * B]
5:  b2done, b2t = [False * B], [0 * B]
6:  while not b2done.all() do
7:     decs, states = decoder(state, predictions)
8:     token-probs = joiner(encs[b2t], decs)
9:     predictions = argmax(token-probs)
10:     b2t[b] += 1 where prediction[b] is Ø
11:     b2done[b] = true where b2t[b] $>$ audio-length[b]
12:     while some prediction is blank and not done do
13:        token-probs = joiner(encs[b2t], decs)
14:        more-predictions = argmax(token-probs)
15:        predictions[b] = more-predictions[b] where prediction[b] is Ø
16:        b2t[b] += 1 where prediction[b] is still Ø
17:        b2done[b] = true where b2t[b] $>$ audio-length[b]
18:     hyps.append(predictions)
19:  return hyps
3.3 Label-looping algorithm for TDT models
Token-and-duration Transducers (TDT) [11] is an extension of Transducer models that decouples token and duration prediction. With TDT, $t$ can be updated regardless of whether the predicted token is blank, and can be incremented by more than 1. The label-looping algorithm for the TDT model is mostly similar to that of conventional Transducers. The only difference is at lines 10 and 16 when updating b2t, the increment amount of the TDT model comes from the model’s duration prediction. Due to space limitations, we refer our readers to our open-source implementation for the actual algorithm.
3.4 Precomputation of encoder/decoder projections
In the standard Transducer model architecture, two linear layers are required to project the encoder and decoder outputs into the same vector space. While this computation is not costly per se, it is repeatedly run through the decoding process and thus can become a significant overhead. In our algorithm, we propose precomputing those projections, at lines 2 and 7 in Algorithm 3, before feeding them to the joiner.
We point out it is also possible to precompute the projections with the original Algorithm 1 (lines 2, 6, and 14). However, the decoder updates (lines 6 and 14) happen much more frequently, and part of the output would be filtered out at lines 7 and 15. As a result, the precomputation of projections would have a smaller impact on the original algorithm.
4 Experiments
We compared our label-looping algorithm against the baseline algorithm shown in Algorithm 2. All experiments are conducted with the NeMo [1] toolkit. Table 1 shows the information of models that we use for evaluation. All these models are publicly available. They use 8X subsampling in its Fast-Conformer [18] encoder’s output, and 1024 BPE [19] vocabulary size at the output side.  [^1] The model names are stt_en_fastconformer_transducer_large, parakeet-rnnt-1.1b, stt_en_fastconformer_tdt_large, parakeet-tdt-1.1b respectively. Modes can be accessed by appending the model name after https://huggingface.co/nvidia/. For each model, we report its inversed real-time factor 222defined as the ratio of audio length to its decoding time. E.g. if it takes 1 second to decode a 2-second audio, then RTFx = 2. (RTFx) on the LibriSpeech [20] test-other dataset. To provide a better picture, we also include the RTFx for time spent in decoding only, excluding encoder computations, since the latter can more directly reflect the speed-up brought by our algorithms. To get a more accurate measurement, we first run decoding twice to “warm up” the cache required for those algorithms to work more efficiently, and report the average time from the third to fifth measurement. All experiments are done on NVIDIA RTX A6000 GPU.
Table 1: Transducer models used for evaluation
model model type # params (millions)
RNNT-L Transducer 114
RNNT-XXL Transducer 1100
TDT-L TDT 114
TDT-XXL TDT 1100
4.1 Results with Transducers
Table 2 presents our results with conventional Transducers. We see the label-looping algorithm consistently speeds up decoding regardless of batch size. Also, a larger relative speedup can be seen for larger batch sizes. This is because, as the batch size grows, the baseline algorithm introduces more overhead, since the more utterances, the more likely one utterance would need to wait for other utterances to advance $t$ in decoding. We also observe greater speed-up for non-encoder computations. Notably, for both models, our algorithm brings around 2.6X speed-up for batch-size=32 on non-encoder computations.
Table 2: Transducer (RNN-T) decoding speed of different algorithms and batch-sizes on LibriSpeech test-other. We report “total computation RTFx / non-encoder RTFx”, where RTF is the inversed real-time factor. Bf16 precision is used for all models. WER is 3.9% for the L model, 2.7% for the XXL model.
model-batch baseline ours rel. speedup
L-1 70.1 / 105.6 95.5 / 179.0 1.4X / 1.7X
L-4 158.4 / 200.1 257.1 / 385.4 1.6X / 1.9X
L-16 313.7 / 390.1 584.5 / 894.4 1.9X / 2.3X
L-32 383.8 / 512.8 751.2 / 1403.6 2.0X / 2.7X
XXL-1 47.2 / 98.8 59.3 / 174.3 1.3X / 1.8X
XXL-4 123.9 / 191.3 183.1 / 380.0 1.5X / 2.0X
XXL-16 218.0 / 388.5 321.6 / 886.2 1.5X / 2.3X
XXL-32 248.4 / 532.7 352.2 / 1393.4 1.4X / 2.6X
4.2 Results with Token-and-Duration Transducers
Table 3: TDT decoding speed of different algorithms with different batchsizes on LibriSpeech test-other. Total RTFx / non-encoder RTFx. WER is 3.7% for the L model, 2.8% for XXL for our label-looping algorithm; for baseline, the WER fluctuates.
model name baseline ours rel. speedup
L-1 116.0 / 264.5 116.5 / 269.3 1.0X / 1.0X
L-4 243.4 / 358.1 347.7 / 632.6 1.4X / 1.8X
L-16 416.2 / 563.9 825.3 / 1615.9 2.0X / 2.9X
L-32 477.2 / 696.7 1006.8 / 2634.2 2.1X / 3.8X
XXL-1 66.1 / 241.3 66.1 / 251.7 1.0X / 1.0X
XXL-4 147.9 / 258.8 219.5 / 584.5 1.5X / 2.3X
XXL-16 243.7 / 478.3 374.8 / 1467.9 1.5X / 3.1X
XXL-32 265.2 / 616.3 394.0 / 2434.1 1.5X / 3.9X
Our TDT model experiments are shown in Table 3. Note, [11] states that TDT’s batched decoding is tricky, and proposes an approximate algorithm, which is not deterministic and requires special modifications during model training. At the time of writing, there is no open-sourced implementation of exact and general-purpose TDT batched decoding, and ours is the first such implementation available to the public. For the baseline, we adopt the approximate method from [11], which takes the minimum of predicted durations in the batch for advancement. Note: this method has non-deterministic outputs and is only included here for time comparison purposes. We can see from the Table that greater speed-ups can be seen with TDT models than conventional Transducers, and we see over 3.8X speed-up for non-encoder computation for both L and XXL models.
5 Analysis
5.1 Impact of precomputing encoder/decoder projections
In this section, we study the effects of precomputing the encoder and decoder projections, where we compare the decoding RTFx with or without the precomputation of projections. The results are in Table 4. We see that overall, precomputing the encoder and decoder outputs brings over 20% speedup for the decoding process excluding encoder computations.
Table 4: Decoding RTFx (excluding encoder) between precomputation of projections and on-the-fly projections. Decoded on LibriSpeech test-other with RNNT-Large, bf16 precision.
batch-size w/o w/precomputation rel. speed-up
1 147.7 179.0 1.21X
4 303.8 385.4 1.27X
32 1080.3 1403.6 1.30X
5.2 Impacts of different subsampling rates
In all experiments we report above, the encoder uses 8X subsampling. Another commonly used subsampling rate is 4X and we test our model’s speed on a publicly available model  [^3] https://catalog.ngc.nvidia.com/orgs/nvidia/teams/nemo/models/stt_en_conformer_transducer_large and show the results in Table 5. We see that with 4X models, our algorithm also significantly improves inference speed, up to 1.8X for total runtime and 2.6X for non-encoder computation.
Table 5: Performance on conformer-large models with 4x subsampling, bf16 precision. We report the total / non-encoder RTFx for decoding Librispeech-testother.
batch-size baseline ours rel. speed-up
1 48.2 / 63.2 79.5 / 130.8 1.6X / 2.1X
4 104.7 / 123.1 192.9 / 262.3 1.8X / 2.1X
16 187.2 / 243.1 341.6 / 577.5 1.8X / 2.4X
32 226.0 / 328.7 399.0 / 866.2 1.8X / 2.6X
5.3 Combining with TorchScript and CUDA Graphs
Since the speedup brought by our method is purely algorithmic, it can be combined with other methods that speed up the decoding process from other aspects such as code compilation and GPU call optimization. Here we report decoding results combining our algorithm with TorchScript [^4] https://pytorch.org/docs/stable/jit.html and with CUDA graphs [21] methods. Note, TorchScript and CUDA graph techniques aren’t compatible at the time of writing, hence we do not include experiments combining them both.
Table 6: Combining Label-looping with TorchScript or CUDA graphs. Batchsize = 32. Full-time RTFx / Non-encoder RTFx.
model RNNT TDT
label-looping 751.2 / 1403.6 1006.8 / 2634.2
+ TorchScript 882.1 / 1942.4 1118.0 / 3628.2
+ cudagraphs 1232.7 / 5197.2 1393.4 / 9614.8
Note, that we show this table to see that our method is compatible with other methods. Because of the scope of this paper, this is the only table with results that combine those other optimizations. The speedup reported from all other tables is solely from the algorithmic changes of the label-looping algorithm.
6 Conclusion and Future Work
In this paper, we introduce a novel set of algorithms to improve the efficiency of Transducer models. The algorithm uses a label-looping design, which minimizes the number of decoder calls, greatly reducing the runtime of decoding. Our experiments show that label-looping algorithms bring consistent speed-up over baseline batched decoding algorithms. In particular, we see a speed-up of 2.0X for batch size = 32 for conventional Transducers, and 2.1X for TDT models. Since our method is purely algorithmic, we show that it can combine with other methods like TorchScript and CUDA graphs to bring further speed up to Transducer decoding. Our algorithm will be open-sourced through the NeMo [1] toolkit.
We will continue our research efforts to improve the algorithm for future work. In particular, since our algorithm reduces the computational cost of the decoder greatly, this allows us to scale up the decoder for Transducer networks, e.g. using more layers or using more sophisticated networks like Transformers.